OpenStudy (anonymous):

open file and help with #2 and #3 please??

OpenStudy (anonymous):

2 and 3??

OpenStudy (anonymous):

@nikato @mary.rojas @SolomonZelman @eliassaab @ranga @francoisderozay @dinnertable @math97 @lonnie455rich @Ostyn17 @caitlynmcdaniel @Calcmathlete

OpenStudy (anonymous):

ANYONE?!

OpenStudy (solomonzelman):

I would suggest to write them so that they are readable so that we can help you.

OpenStudy (shamil98):

Or take a better picture.

OpenStudy (anonymous):

just zoom in if needed? yes? or no option for that?

OpenStudy (shamil98):

You are going to help your grandmother create a garden of tullips and daffodils. She plans to plant more daffodils than tullips .The florist tells you tullips cost $.50 per bulb and $.75 bulb for daffodils. How many of each can you purchase if your budget is $52?

OpenStudy (shamil98):

There, that's number two.

OpenStudy (anonymous):

That is number 1^^ i need number 2 and 3 aha..

OpenStudy (shamil98):

No, that's number two.

OpenStudy (shamil98):

It's labeled as 2)

OpenStudy (nikato):



OpenStudy (anonymous):

fuuck im sorry, i ment 3 and 4?!

OpenStudy (shamil98):

You should specify which one for future references..

OpenStudy (anonymous):

i forgot my teacer numbered them wrong!

OpenStudy (nikato):

for 3, let x be the number of pounds of coffee A and 6-x=coffee B

OpenStudy (anonymous):

so what are the equations?

OpenStudy (nikato):



OpenStudy (nikato):

it'll the the 3rd colomn. the first row plus the second row = 3rd row

OpenStudy (anonymous):

so it is x(8.4)+(6-x)7.28=6(7.95) ?? @nikato

OpenStudy (anonymous):

and then solve? lol

OpenStudy (nikato):

yea, then just solve for x. and remeber 6-x is what u r looking for

OpenStudy (anonymous):

what do you mean im looking for 6-x?

OpenStudy (nikato):

becuz ur question is asking for how many pounds of the 7.28 coffee. and based on that graph, that coffe is 6-x pounds

OpenStudy (anonymous):

so i got an asnwer, can you do it too just to make sure? im sorta iffy about mine!

OpenStudy (anonymous):

i will give you a badge @nikato and an award :D haha

OpenStudy (nikato):

ok, so what is ur answer?

OpenStudy (anonymous):

someone please solve #4

OpenStudy (anonymous):

i got 1.2496